OpenStudy (anonymous):

could someone teach me how to solve this? Solve the equation. show your work. check your answer. 3x - 7 =5x + 19

OpenStudy (anonymous):

can you help us again? @FibonacciChick666

OpenStudy (anonymous):

Yeah sure. see subtract 3x on both the sides of the '='

OpenStudy (anonymous):

Add 7 on both sides -_-

OpenStudy (fibonaccichick666):

ok, so the goal here is to have only x terms on one side, and only non x terms on the other

OpenStudy (fibonaccichick666):

so first, can you tell me what the x terms are?

OpenStudy (anonymous):

3x and 5x

OpenStudy (anonymous):

ill make this a lot easier on you

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (fibonaccichick666):

good, so now, what are the non-x terms

OpenStudy (anonymous):

enter the equation onto that

OpenStudy (fibonaccichick666):

@skate-4-life270, 1. we are here to help not give answers. openstudy.com/code-of-conduct 2. some people actually want to learn. 3. please bug off if that is what you call a meaningful contribution.

OpenStudy (fibonaccichick666):

sorry, @AngelofVengeance6661, what are the non-x terms again?

OpenStudy (anonymous):

i didnt give her the answer, that site shows her step by step how to do it

OpenStudy (anonymous):

and i know about he code of conduct youve told me it 100 time the past week, and what i did isnt violating it

OpenStudy (fibonaccichick666):

that yields a step by step solution with no resoning nor assistance hence, just an answer

OpenStudy (anonymous):

the non-x terms are -7 and 19

OpenStudy (fibonaccichick666):

good, so now, pick a side. Which side do you want to put the x terms on?

OpenStudy (anonymous):

um. right?

OpenStudy (fibonaccichick666):

ok so we shall have non x terms on the left and x terms on the right

OpenStudy (fibonaccichick666):

\[3x-7=5x+9\]

OpenStudy (fibonaccichick666):

the first thing we want to do is figure out how to 'undo' the addition of the 9 on the right

OpenStudy (fibonaccichick666):

so what is the opposite of addition?

OpenStudy (anonymous):

-

OpenStudy (fibonaccichick666):

yep subtraction, so to undo the addition of 9, what do you think we need to do to both sides?

OpenStudy (anonymous):

subtract

OpenStudy (fibonaccichick666):

subtract what?

OpenStudy (anonymous):

wait a minute im lost o.o

OpenStudy (fibonaccichick666):

ok so just reread the last two comments I have made, think "what am I undoing?" "How do I undo it?"

OpenStudy (anonymous):

wait do you do this the same way you do a regular problem just with two numbers instead of 1

OpenStudy (fibonaccichick666):

I think that what your saying is true, but I don;t know what you are saying

OpenStudy (fibonaccichick666):

so, give me an example of what you are saying so I can see if it is

OpenStudy (anonymous):

i understand it now. i even checked it myself c:

OpenStudy (fibonaccichick666):

ok :) then show me your work

OpenStudy (fibonaccichick666):

try and explain every step, you'll understand better if you do

OpenStudy (anonymous):

3x-7=5x-9 -5x from both sides = 2x-7 and -9 +7 both sides =-2x and -2 divide by -2 x=1

OpenStudy (some.random.cool.kid):

yeah if your going to use an online source it has to explain the math and it has to be accurate and no wiki.

OpenStudy (fibonaccichick666):

ah so what is 3-5?

OpenStudy (fibonaccichick666):

ok wait, you may have it you corrected it in the next step

OpenStudy (fibonaccichick666):

so here is how I would write it, let's see if we get the same:\[3x-7=5x\color\red{+19}\]

OpenStudy (anonymous):

ur way takes too long

OpenStudy (fibonaccichick666):

so our first thing is we didn't copy the problem down correctly, but let's see if you did the steps after correct.

OpenStudy (fibonaccichick666):

@skate-4-life270 I help and teach. That is what this site is for. DOn't like it? then leave

OpenStudy (anonymous):

ok i will

OpenStudy (fibonaccichick666):

ok so you wrote,\[3x-7=5x-9\] subtract 5x from both sides-I agree\[3x-5x-7=5x-5x-9\]\[-2x-7=-9\] Add 7 to both sides, (I agree with the step) \[-2x-7+7=-9+7\]\[-2x=-2\] Divide by -2. agreed \[\frac{-2x}{-2}=\frac{-2}{-2}\] \[x=1\]

OpenStudy (fibonaccichick666):

so end result, you did the math correctly, but if you do not copy the problem down correctly, you will unfortunately be incorrect